Instrument Log 

25 MT T/H "B" (West) Side       C1   Lamont, OK

Date

Operator

Comments

11/16/2015

MS

Carriage was lowered for work on 60m CO2 flux.  Data for this instrument may be affected around 1815 GMT to 1830 GMT.

10/26/15

CAM

D. Cook

Instrument mentor on site for annual 25 and 60M T/H calibration change outs.  Lowered the West booms at 1454 GMT for maintenance.   R/R 25M *T/H sensor.  Raised booms and RTS at 2012 GMT. 

*Old – S2920007   *New – S2920009

7/20/15

PAD/JEM

West side booms were lowered to remove 25 and 60M IRGA and R/R 25 and 60M Sonic Anemometers.  Booms were lowered at 1405 GMT, raised at 2050 GMT.  No issues known with this instrument.

7/6/15

JEM

There was a power outage overnight and the TWR25M did not resume normal operation after the power was restored.  Found no program loaded on the datalogger.  R/R dataloggers internal lithium battery.  Reloaded the program @1517 GMT.  Verified data looked normal and RTS @1527 GMT.

4/2/15

PD

25M CO2 Flux program was stopped on 4/1/15 so computer could be relocated to temporary EF enclosure at the base of the 60M tower.  It failed to restart correctly.  Powercycled side B, 25M instruments at 1445 GMT.  RTS at 1500 GMT.  No issues known with this instrument.

3/13/15

JEM

Side B of tower was dropped to remove IRGAs for calibration on 3/12/15.  IRGAs need to be installed.  Lowered tower at 2003 GMT, installed IRGAs and raised tower.  RTS at 2123 GMT.

3/12/13

JEM

Side B of tower was dropped to remove IRGAs for calibration.  Lowered tower at 1752 GMT, removed IRGAs and raised tower.  RTS at 1842 GMT.

1/27/15

JEM

The boom was lowered for MFR25M maintenance.  Data will be effected from 2105 GMT to 2120 GMT.

10/28/14

CAM

D. Cook

Instrument mentor on site for annual 25 and 60M T/H calibration change outs.  Lowered the West booms at 1506 GMT for maintenance.   R/R 25M *T/H sensor.   Performed ice bath calibrations on PRTD sensor.  Raised booms and RTS at 2230 GMT. 

*Old – S2920008   *New – S2920007

6/23/14

CAM

No data had been collected from 25 T/H since 1205 GMT on 6/21.  Found the UPS that powers the comms equipment in the AOS trailer was down.  Power cycled UPS and restored power at 1220 GMT.  Normal data collections resumed at next interval. 

5/27/14

CAM

No data had been collected from instrument since 1208 GMT on 5/25.  Found the UPS that powers the comms equipment in the AOS trailer was down.  Breaker for UPS supply line was tripped.  Reset breaker at 1325 GMT.  UPS powered up and comms equipment power restored.  Normal data collections resumed at next interval. 

11/14/13

JEM

West tower booms were lowered to work on the 25M CO2FLUX system.  Any 25M T/H data between 1600 GMT and 2300 GMT should be ignored.  RTS @2300 GMT.

6/17/13

JEM

The West booms need to be lowered to reinstall the 25M and 60M IRGA.  25M T/H data is affected between 1947 GMT and 2135 GMT.  RTS @2135 GMT.

6/13/13

JEM

West booms were lowered from 1455 GMT to 1630 GMT so the 25M and 60M IRGA’s could be calibrated.  25M T/H data is affected during this time.  RTS @1630 GMT.

4/24/13

Twr Sys

Tower Systems on site for tower maintenance.  Unsnagged SW winch cable from foot pegs and straightened bent sections of boom guide rails.  Adjusted boom vertically on winch cable for proper alignment with top connectors.  Winch cable will need to be remarked next time the booms are completely lowered.

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.  Power was restored at 1920 GMT.  Instrument RTS status O.K and data collections resumed once SDS operations were restored.

2/14/13

PAD/KT

Issue with 25M CO2 flux instruments.  Mentor requested 25M side B of tower power cycled to see if that would correct the issue.  Power cycled 25M side B of tower.  No issues known with this instrument.  RTS at 2215 GMT.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.  Power was restored at 2115 GMT after electrical upgrades were complete. 

11/13/12

CAM

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound.  Power was restored at approximately 1930 GMT after all transformers were replaced.  Tower instruments resumed normal operation.

10/11/12

JEM

The tower needs to be lowered to reinstall the 25M and 60M IRGA's.  Lowered the booms @1753 GMT.  The booms were raised and RTS @1915 GMT.

10/10/12

CAM

D. Cook

The west 60M temperature sensor has been reading erratically since the calibration was done yesterday.  Lowered west booms at 1905 GMT.  T/S found one of the sensor wires was pinched against the ventilator housing and intermittently shorting out.  Repaired wire and tested ok.  Raised boms and RTS at 2045 GMT.

10/9/12

CAM

D. Cook

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts.  Lowered the West booms at 1845 GMT for maintenance.   R/R 25M *T/H sensor.   Raised booms and RTS at 2210 GMT. 

*Old – S2920007   *New – S2920008

3/8/12

PAD

Radon Gas Analyzer inlet pipe has separated approximately 40M above the ground.  Contractor is on site today to climb tower and repair inlet tube.  Power to tower needs to be shut down before contractor can climb.  Shut down power at base of tower to 25 and 60M levels, side A at 2023 GMT.  After repairs were made and complete inspection of inlet pipe was performed, power was restored.  No issues with this instrument are known at this time.  RTS at 2257 GMT.

2/14/12

PAD/CAM

25M IRGA was calibrated and needs to be installed on tower.  Lowered side B of tower at 2023 GMT.  Raised tower at 2105 GMT.  No issues known with instrument at this time.

2/7/12

PAD

25M IRGA needs to be calibrated.  Lowered side B booms at 1414 GMT.  Removed IRGA and raised tower at 1500 GMT.  No issues known with this instrument.

11/1/11

CAM

D. Cook

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts.  Lowered the West booms at 1758 GMT for maintenance.   R/R 25M *T/H sensor.  Performed ice bath calibrations on PRTD sensor. Raised booms and RTS at 2142 GMT. 

*Old - S2920011  *New – S2920007

4/26/11

PAD

Calibrated 25 and 60M IRGAs and they need to be installed on tower.  Lowered side B of tower at 1721 GMT.  Raised at 1800 GMT.  No problems known with instrument at this time. 

4/26/11

PAD

Mentor requests 25m IRGA calibrated.  Lowered west boom of tower at 1354 GMT.  Raised tower at 1434 GMT.  No problems known with instrument at this time.

4/15/11

JEM

The instrument has not collected since the 0400 hour.  Due to other instruments that are down the source of the problems seems to be the Rocketport hub.  The Rocketport power light is flashing.  Power cycling the Rocketport hub did not correct the issue.  Notified the IT department of the issue.

10/19/10

CAM

Tower Systems on site for tower inspections.  Need to secure power for maintenance.  Secured AC power to 25m and 60m levels at 1407 GMT.  Restored power following inspection at 1620 GMT.

10/18/10

CAM

D. Cook

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts.   Lowered the West booms at 1423 GMT for maintenance.   R/R 25M *T/H sensor.   Performed ice bath calibrations on PRTD sensor. Raised booms and RTS at 2026 GMT. 

*Old - S4930004  *New – S2920011

11/19/09

CAM

SDS instrument status page indicated no data had been collected since 2108 GMT on 11/18.  T/S found the UPS in the AOS trailer which powers the tower communications equipment had no output.  AC LINE and ALARM LEDs were lit on the UPS.  Power cycled UPS at 1342 GMT and power was restored.  RTS at 1345 GMT. 

9/21/09

D. Cook

CAM

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts.   Lowered the West booms at 1506 GMT for maintenance.   R/R 25M *T/H sensor.   Performed ice bath calibrations on PRTD sensor. Raised booms and RTS at 1750 GMT. 

*Old - S1720002 (HMP35D)   *New – S4930004 (HMP45D)  

7/1/09

PAD

25M IRGA needs calibrated.  Lowered tower at 1424 GMT to removed IRGA.  Tower was raised and 25M T/H RTS at 1604 GMT.  No problems know with instrument at this time.

12/22/08

CAM/PAD

Power was removed from 25M T/H datalogger on 12/16 due to a dead short on the 12V sensor wire at the 25M level.  The short was killing the power supply common to both the 25M and 60M dataloggers.  The booms could not be lowered at that time due to freezing precipitation so the 25M datalogger was disconnected to allow 60M T/H to continue to operate.  Ice has now thawed.  Lowered the booms at 1955 GMT to T/S short at 25M level.  Found a piece of aluminum shaving had fallen into the DC connector on the 25M boom when the 60M boom was drilled to mount a camera on 11/18.  The aluminum shaving was shorting the 12V power pin to ground.  Removed shaving and restored power to 25M datalogger.  Raised booms and RTS at 2029 GMT.  Confirmed good power and sensor readings.  RTS at 2035 GMT.

12/16/08

PAD

Last SDS collection was on 12/11/08 at 2308 GMT.  West side 25M and 60M dataloggers share a common power supply.   Chris Martin and I found low voltage at dataloggers. Suspect short in tower connector at 25M level or possibly T/H sensor.  Cannot lower tower due to ice, but will further troubleshoot when weather conditions allow.  Disconnected 25M datalogger and sensor from power supply to allow 60M instrument to operate.  RTS at 1745 GMT.

12/15/08

PAD

DQPR 2079 states:  "This problem is for the sgp1twr25mC1.b1 datastream. The rh values are currently failing the MIN check and temperature, vapor pressure and battery voltage are missing".  Mentor requested West side tower booms lowered and raised to reseat connections.  This was done on 11/18/08.  SDS has not collected data from 25M and 60M T/H sensors since that date.  T/S found low voltage at 25 and 60M T/H dataloggers.  When 25M sensor is disconnected from datalogger, voltage at dataloggers returns to 13.6V.  Lowered booms and checked ventilator fan and sensor.  No problems found.  Raised tower and verified we could communicate with dataloggers.  RTS at 1630 GMT.

11/18/08

PAD

Facilities department needs to lower West side 25 and 60M booms to install camera on tower.  Tower was raised at 1725 GMT.  No problems with this instrument.

11/13/08

PAD

Facilities department needs to lower W side booms to mount tower camera.  Lowered tower at 1450 GMT.  Raised tower at 1537 GMT.  No problems with this instrument known.  RTS at 1540 GMT.

5/29/08

PAD

25M IRGA needs to be installed after calibration.  Lowered tower at 1926 GMT and installed IRGA.  Raised tower at 2006 GMT.  No problems known with this instrument at this time.  RTS at 2006 GMT.

5/28/08

PAD

Mentor requests 25 and 60m IRGAs located on side B, West side of boom calibrated.  Lowered side B, West side 25m boom at 1529 GMT and removed IRGA.  Raised boom at 1950 GMT.  No problems with this instrument known at this time.

4/21/08

PAD

IRGA SN 75B/H-0435 was returned to the CF from LiCor after repairs.  Mentor request unit placed on 25m level of 60m tower and RTS.  At1820 GMT, Mark Klassen and I lowered west booms, side B, of 60m tower.  Booms were raised and instrument RTS at 1850 GMT.  No problems known with this instrument.

2/28/08

PAD

Mentor request 25M IRGA removed from tower.  Lowered tower at 1853 GMT.  Raised tower and instrument RTS at 1923 GMT.  No problems known with this instrument at this time.

11/1/07

PAD

25M boom on West side was power cycled to reset IRGA.  No problems with instrument at this time. 

10/3/07

PAD

25M IRGA was removed for calibration and needs to be RTS.  Lowered West side B of tower to install instrument.  No problems with 25M T/H known at this time.  RTS at 1806 GMT.

10/1/07

PAD

25M IRGA needs calibrated.  Lowered West side B of tower to remove instrument.  No problems with 25M T/H known at this time.  RTS at 1806 GMT.

7/26/07

D. Cook

CAM

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts.   Lowered the West booms at 1354 GMT for maintenance.   R/R 25M *T/H sensor.   Performed ice bath calibrations on PRTD sensor. Raised booms and RTS at 1900 GMT. 

*Old – S2920010 (HMP35D)   *New - S1720002 (HMP35D)

4/5/07

PAD

Side B of tower needs to be lowered to install 25M IRGA.  Tower was lowered at 1537 GMT and 25M level was power cycled troubleshooting power issue with 25M IRGA.  No problems with instrument at this time.  RTS at 1930 GMT.

4/2/07

PAD

West booms, side B, needs to be lowered to remove 25m IRGA.  Let down side B booms at 1428 GMT.  Boom were raised and RTS at 1503 GMT.  No problems with instrument at this time. 

10/11/06

D. Cook

CAM

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts.   Lowered the West booms at 1505 GMT for maintenance.   R/R 25M *T/H sensor.   Performed ice bath calibrations on PRTD sensor. Raised booms and RTS at 2117 GMT.  *Old – S1720002 (HMP35D)    *New – S2920010 (HMP35D) 

10/5/06

PAD

25M IRGA needs to be calibrated.  Lowered side B (West side boom) at 1745 GMT and calibrated 25M IRGA.  No problems with this instrument.  RTS at 2035 GMT.

5/24/06

MK

The tower is being lowered for maintenance on the 25M and 60M  LI-7500 units. Lowered tower at 1328 GMT, raised tower at 1630 GMT.

5/22/06

MK

The tower is being lowered for maintenance on the 25M and 60M  LI-7500 units. Lowered tower at 2001 GMT, raised tower at 2040 GMT.

5/18/06

PAD

25M IRGA on tower is scheduled for calibration.  Lowered 25M side B of tower at 1526 GMT.  Removed 25m IRGA for calibration.  Raised 25M boom at 2102 GMT.  No problems with this instrument at this time.

12/5/05

MK

The OS needs to be upgraded on all CR10X's. Downloaded current data off of data logger, retrieved program from data logger, loaded new OS.  Loaded program back into CR10X datalogger. System is operating normal. S/N 23756. RTS at 2055 GMT.

9/6/05

PAD

Power was shut off to PGS shed to make repairs to conduit between PGS shed and tower.  PGS shed instruments were restarted at 1825 GMT.  Could not get PGS computer to talk to 25m Sonic and IRGA.  Power cycled side B 25m instruments on tower twice between 1840 GMT and 1900 GMT.  No problems with this instrument at this time.  RTS at 1915 GMT.

8/5/05

D. Cook

CAM

Instrument mentor back on site to finish the annual 25 and 60M T/H sensor changeouts and calibration checks.  Lowered west booms at 1329 GMT.    R/R 25M *T/H sensor.  Performed ice bath calibrations on temperature sensor.  Raised booms and RTS at 1647 GMT.   *Old – S2920008 (HMP35D)  *New – S1720002 (HMP35D)

8/4/05

D. Cook

GRS

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts and temperature sensor ice baths.  Tower System techs also on site for tower inspection.  Turned the 25 and 60M levels power off at 1330 GMT so Tower Systems could service tower. Lowered the west booms at 1330 GMT for inspection.  Made power and data connections for the 25M boom at ground level, but did not bring the 60M boom to ground level.  Had to cut the maintenance short and raise the booms due to lightning in the area.  Raised west booms and RTS at 2001 GMT.

7/29/05

PAD

Instrument mentor request 60M IRGA placed in service.  No problems with instrument at this time.  Shut down 25M Sonic 2 program at 1843 GMT.  Lowered West booms and installed 60M IRGA, SN 0435.   Raised tower and RTS at 2013 GMT.

2/22/05

PAD

Need to power cycle other instruments on tower to reestablish communication with computer.  Power cycled instruments at 25m level on west side of boom at 1752 GMT.  RTS at 1807 GMT.

10/20/04

MR

The 25M elevator needs to be lowered due to maintenance on other equipment.  Lowered the elevator at 1750 GMT.  Raised the elevator at 1930.  RTS at 1935.

10/13/04

MR

Maintenance on other equipment has been completed and the tower needs to be raised.  Raised the tower at 1630 GMT.  RTS at 1630 GMT.

10/12/04

MR

Lowered the tower at 1416 for IRGA maintenance. Removed the 25 Meter and 60 Meter IRGAS for calibration.  Both 25M and 60M IRGA's need to be calibrated and the internal chemicals changed. Will leave the tower down overnight while these instruments are being calibrated.

9/30/04

JEM

Collected data from the datalogger @1947 GMT due to a collection problem with the SDS.  RTS @2000 GMT.

9/27/04

D. Cook

CAM

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts and temperature sensor ice baths.  Lowered west booms at 1432 GMT.    R/R 25M *T/H sensor.  Performed ice bath calibrations on temperature sensor.  Raised booms and RTS at 2105 GMT.   Encountered some difficulty getting the 25m signal connector to mate when raising booms due to the upper connector being tilted slightly.  Mentor recommends not lowering the booms unless absolutely necessary until Tower Systems can come out to inspect/repair the connector.   *Old – S1720004  (HMP35D)   *New – S2920008 (HMP35D)

8/30/04

MR

The instrument mentor wants the SW elevator bumped so the connections are reseated.  Bumped the SW 60M elevator.  RTS at 1830 GMT.

8/5/04

MR

The SW elevator needs to be lowered due to SN# need to be checked.  Lowered the tower at 1450 GMT.  Raised the tower at 1523.  RTS at 1530 GMT.

7/27/04

MR

This is for the SW T/H Probe.  The instrument mentor has noticed some data missing between 6/30/04 @ 16:30 and 7/1/04 @ 15:15 GMT and wants to see if it can be recovered from the storage module. Downloaded all the data from the storage module as per mentors instructions.  RTS at 1910 GMT.

6/3/04

PAD

During troubleshooting 4m CO2flux system network problems, needed to power cycle 25M W boom.  Power cycled 25M west boom at 2030 GMT.  There is no failure with this instrument.

4/29/04

MR

Work on other equipment has been completed and the west 25M tower can be raised.  Raised the 25M tower.  RTS at 1341 GMT.

4/26/04

JEM

The IRGA’s (LI-7500) need to be removed so they will be ready for Mark Fisher to calibrate on 4/27/04.  Closed the program @1527 GMT.  Removed the 4M IRGA @1610 GMT.   Lowered the 25M boom to the ground.  Removed the 25M IRGA @1535 GMT.  Lowered the 60m boom to the ground.  Removed the 60M IRGA @1547 GMT.  Placed components in the GIF trailer for Mark Fisher.  Left booms at ground level pending calibration of the IRGA’s.

2/11/04

MR

The elevators need to be lowered due to work on other equipment. Lowered the elevators at 1614 GMT.  Raised the elevators at 2230 GMT.

2/10/04

MR

Tower systems is finished with the tower and the elevators need to be returned to their operating position.  Raised the towers at 1800 GMT.  RTS at 1830 GMT.

2/9/04

MR

Tower systems is here and will be working on the 60M tower connections. Lowered the 25M and 60M elevators at 1600 GMT.  They will remain lowered until tower systems is finished with the work.

1/29/04

PAD

Instruments on 25M West boom were power cycled due to problems with 60M CO2 flux.  Power cycled 25M instruments at 2209 GMT.  RTS at 2224 GMT.  No problems with this instrument.

10/6/03

MR

The instrument mentor wants the internal battery checked on the data logger for the 25M SW sensors.  Checked the battery.  The reading is 3.29 volts.  RTS at 1740 GMT.

8/14/03-8/15/03

CAM

Dave Cook on site for annual 25 and 60M T/H calibration changeouts and temperature sensor ice baths.  Tower Systems also on site for tower inspection and upgrades to PGS filter systems.  West booms lowered at 1515 GMT on 8/14 for Tower Systems maintenance.  The 60M level plungers connectors were discovered to have cracks developing on the boom portion of the connector.  To minimize stress, the booms were left down overnight.  Returned on 8/15 for sensor calibration changes.  R/R 25M *T/H sensor.  Raised booms and RTS at 1939 GMT on 8/15.   *Old – S4520023  *New – S1720004

5/15/03

PAD

25 T/H Side B was let down to perform maintenance on other instruments.  Put up @ 2106 GMT.  RTS @ 2140 GMT. No action performed on this instrument.

3/18/03

JEM

The 25M T/H has not collected since 3/17/03 @1502 GMT.  The power strip that powers the bit driver in the AOS trailer was found powered off.  Powered the power strip back on @1505 GMT.  The power strip was likely powered off by accident while rearranging the AOS system.  RTS @1515 GMT.

2/27/03

MR

The 25M T/H will not communicate with the SDS system.  T/S found a bad fiber optic cable.  R/R the cable and the 60 M T/H cables as well.  Communications was restored.  RTS at 1730 GMT.

2/26/03

MR

The 60M T/H probe that was installed on Friday is not working and is causing the 25M T/H to fail collections.  Lowered the tower at 1658 GMT.  T/S found that the battery for the data logger was dead and that the 60M T/H probe had not been wired properly.  Wired the probe properly and installed a new battery.  The probe began to work and was reading ok by the time the tower was raised.  Notified the mentor of the problem.  Raised the tower at 2022 GMT.

2/21/03

MR

The west side (side B) tower needs to be lowered due to maintenance being performed on other instruments.  Lowered the tower at 1927.  Raised the tower at 2120 GMT.  RTS at 2127 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power  was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

11/1/02

CAM

SDS quicklook plots indicated that the west 25M temperature, humidity and vapor pressure were all giving bad readings since approximately 2200 GMT on 10/31.  T/S found that the 25M boom was not making good contact, as evident by the 25M data line indicator light not being lit on the elevator control panel.  Bumped the west side winch slightly at 1745 GMT and the indicator lit immediately.  Monitored the CR10 from ground level and observed the readings had returned to normal.  RTS status O.K. at 1750 GMT.

8/2/02

D. Cook

Booms lowered from 1354 to 1912 GMT for tower maintenance and 25/60m T/H calibrations.  R/R *T/H sensor for annual calibration.  *Old - S1720002  *New – S4520023

8/1/02

D. Cook

TSI on site.  Booms lowered from 1300 to 2130 GMT for tower maintenance. 

5/14/02

MR

The instrument data needs to be collected and placed on the door step for the SDS to collect.  Collected the data from 1441-1453 GMT.  RTS at 1456 GMT.

8/8/01

MR

The instrument mentor is on site to replace and calibrate the 25M and 60M T/H probes. Lowered the tower at 1356 GMT.  R/R the *25M T/H probe.   Tested and calibrated the probe.  The probe tested good.  Raised the tower at 1849 GMT.  RTS at 1856.   *Old - S4520022   *New - S1720002

2/1/2001

MR/KT

Lowered the SW (B) side of the 60M tower from 2140 GMT to 2125 GMT to secure a loose hygrometer cable at the 60M level. All instruments on the SW side of the tower were affected during this time.

10/30/00

CAM

Instrument mentor sent a new CR10X data logger to replace the old 21X data logger and SM716 storage module on the west side. Waited for the 1805 GMT collection period to complete and then powered the data logger off at 1807 GMT. R/R the 21X *data logger with the new CR10X data logger. R/R the **power supply. The power supply will power both the 25 and 60M loggers. Removed the ***storage module from service. Made all the necessary connections and powered the data logger on at 1930 GMT. Checked all readings and found that the temperature and vapor pressure were reading 99999. Confirmed the program and all connections were correct. Contacted the mentor who advised that the excitation voltage may need to be lowered with the new data loggers. Changed the excitation voltage in the program (table 1, step 6) from 90 to 40 mV at 2110 GMT. The temperature and vapor pressure readings then began to report accurately. RTS status O.K. The old components will be returned to David Cook. *Old - 6948 (21X) *New - 23756 (CR10X) **Old - n/a **New - 19984 ***Old - 5080

10/20/00

CAM

SDS comlog reports indicated that the 25M T/H had not successfully collected files since 2004 GMT on 10/18. Attempted to telnet into the tower COS but was unable to link to either the production or development ports. T/S found that the power supply which powers ports 0 and 4 bit drivers was unplugged. Plugged the power supply in at 1440 GMT and was then able to telnet into both ports. The following data collection was successful at 1505 GMT. RTS status O.K.

8/14/2000

CAM/DC

David Cook on site for annual sensor calibrations and change outs. Lowered all tower booms at 1447 GMT. Connected power and data jumper cables and let sensors stabilize at ground level for 3 hours and checked sensor performance. R/R the west 25M *humidity sensor for annual calibration. Performed ice bath calibration on the temperature sensor. Raised the booms and RTS at 2138 GMT. *Old - S4930005 (model HMP45D) *New - S4520022 (model HMP45D)

11/16/99

CM/SS

Lowered the B side booms from 1610 to 1700 GMT to T/S 60M aspirator fan.

5/19/99

5/20/99

CAM/DC

David Cook on site for annual sensor calibrations. Lowered the 25M B side boom at 1755 GMT on 5/19. Performed ice bath calibration on temperature sensor. Sensor checked out O.K. R/R the *humidity sensor for calibration. New sensor checked out O.K. Left the boom at ground level overnight for ECOR maintenance. Raised B side boom and RTS status O.K. at 1755 GMT on 5/20/99. *Old - S4930004 (Model HMP45D) *New - S4930005 (Model HMP45D).

7/16/98

7/17/98

CAM/DC

David Cook on site for 6 month calibrations. Lowered the booms at 1458 GMT on 7/16/98. R/R *Humidity sensor for 6 month calibration. Performed ice bath calibration on temperature sensor. Left booms connected at ground level overnight. Returned on 7/17/98 and rechecked sensors. Status O.K. Raised booms and RTS at 1513 GMT. *Old - S4930005 *New - S4930004

6/23/98

MW

SDS operations reported a loss of communications with all of the 60M tower instruments beginning at 1004 GMT today. Power cycled the tower COS at 1404 GMT. Communications restored after the power cycle. RTS at 1405 GMT.

2/13/98

CAM/MR

Mentor indicated that the 25M B humidity sensor was reporting near zero constantly. Lowered 25M B boom at 1930 GMT. *R/R humidity sensor and verified proper reading. Raised boom and RTS at 2007 GMT. * Old – S1720002 (Model HMP35D). * New S4930005 (Model HMP45D).

1/7/98

CAM

SDS TR indicated that no files had been received from the 25M T/H side 'B' since 1304 GMT on 1/6/98. T/S on 1/7/98 found that the 'A' side 25M GFCI breaker had tripped on 1/4/98 at 0935 GMT. The data logger power supply for the 25M T/H was plugged into this outlet due to recent problems with the 'B' side 25M GFCI caused by faulty ECOR Sonic electronics. The Sonic electronics had since been replaced so the 25M T/H power supply was moved back to its proper location at 2035 GMT on 1/7/98 and power was restored. Recompiled data logger program and RTS at 2040 GMT. Note that since the breaker initially tripped on 1/4 the data logger had been running on backup battery power until it finally died on 1/6. Data during that period may or may not be affected.

12/19/97

CAM

Lowered booms from 2100 GMT to 2210 GMT for ECOR maintenance.

11/5/97

CAM

While on site for 25/60m humidity sensor replacement, instrument mentor also performed ice bath calibrations on the 'B' (West) side temperature sensors. Both the 25m and 60m booms were lowered on 11/4/97 at 2020 GMT and left connected at ground level overnight. Returned to tower on 11/5 and performed ice bath calibration on the 25m temperature sensor from 1950 to 2010 GMT. Raised booms and RTS at 2135 GMT.

11/4/97

MW

David Cook, instrument mentor, was on site 11/4/97 to replace the 'B' (West)side 25m and 60m Qualimetrics humidity sensors with new Vaisala humidity sensors. Lowered the 'B' side booms at 2020 GMT on 11/4. Connected power and signal jumpers at ground level at 2034 GMT. *R/R the 25m Qualimetrics humidity sensor with new Vaisala humidity sensor. Sensors were left connected at ground level overnight. * Old - model 5134-E, SN 109. * New - model HMP35D, SN S1720002.

10/29/97

CAM

SDS trouble report indicates that although the 25M T/H communications were restored on 10/28, the data files that have been received since then are jumbled. In addition two of the collections were missed completely. Attempted to power cycle the 25M T/H data logger at 1558 GMT on 10/29 but the data logger remained on with the toggle switch in the off position. Disconnected the power supply from the data logger but it still remained on. The transmit light on the bit driver at the data logger end was constantly flashing as was the receive light on the bit driver at the COS end. The bit driver at the data logger end was being powered by both the data logger (via the 'Eagan' adapter) and a bit driver power supply. This arrangement was somehow allowing the data logger to receive power through its 12V output and thus remain on. Unplugged the power supply from the bit driver and the data logger then powered off and the bit driver lights quit flashing. Attempts to monitor the data logger however from the aerosol trailer through the fibers and bit drivers were unsuccessful. R/R both bit drivers and communications were then established at 1725 GMT. SDS data collection was then successful with a good file at 1809 GMT. RTS status O.K.

10/28/97

CAM

SDS trouble report indicated that no files were being received from 25M T/H since 0300 GMT on 10/27/97. T/S on 10/28 found that the "B lower" GFCI circuit breaker at the base of the tower was tripped. The 25M T/H data logger power supply was plugged into this outlet. After the breaker tripped, the 25M data logger would have run on battery power until they drained. Reset the GFCI breaker at 1640 GMT but it immediately tripped again. Suspect that the 25M ECOR Sonic electronics is bad and thus causing the breaker to trip. Moved the 25M T/H data logger power supply to another outlet and RTS at 1645 GMT. 25M T/H communications then resumed at the next collection interval. Turned off the AC power to the 25M "b" side until the boom can be lowered to remove the 25M ECOR Sonic electronics.

9/25/97

CAM

David Cook requested Site Ops verify the data logger parameters and correct if necessary. Checked at 1955 GMT. All parameters were set correctly.

9/04/97

CAM

SDS operator indicated that no files were being received from 25M T/H. T/S found that the AC circuit breaker had tripped on the outlet at the base of the tower. The 25M T/H data logger power supply was plugged into this outlet. After the outlet tripped the data logger continued to run until its batteries were drained at which point it crashed. Reset circuit breaker at 1850 GMT. The data logger program did not start automatically and all values were reporting 0000. Power cycled the data logger at 1920 GMT. This time the program successfully loaded. Data transfer to SDS resumed on the 2005 GMT collection.

8/13/97

GRS

*R/R humidity sensor on the "B" west side at 1830 GMT. Performed calibration checks. Raised tower B on 8/14/97 at 1602 GMT, RTS. * Old - 234 * New - 109.

8/5/97

MW

The 25M boom was lowered at 1803 today in order to service the 25M ECOR instruments. While the boom was down, a stress fracture was found in the aluminum plate where the 25M boom is connected to the elevator cable via a u-bolt. Due to the possibility of a safety hazard, the 25M boom was left in the lowered position. Therefore, the 25M temp/hum sensor is down until further notice.

5/28/97

CAM

David Cook requested Site Ops check the possibility that the 25M and 60M T/H communications lines have somehow been reversed. Disconnected the 25M data logger and had SDS operator attempt to tip the logger. The tip was successful to the 25M but unable to tip 60M thus verifying that the communications were swapped somewhere along the line. Swapped the communications lines at the data loggers as a temporary fix at 2110 GMT. Will research the situation and find where the lines were actually reversed. Meanwhile communications are now correct.

5/21/97

MW

Dave Cook requested the 25M and 60M temp/hum sensors and aspirators be lowered in order to check their operation. Lowered the booms at 1500 GMT. Connected power jumpers and observed both aspirators working properly. Connected the signal lines and monitored the temp/hum sensors after a 10 minute warm up period. The results were as follows: 60M temp = 20.67, 60M hum = 53.37, 25M temp = 20.74, 25M hum = 54.22. Attempted to call Dave Cook while sensors were down but could not get in touch with him. Raised both tower booms and RTS at 1600 GMT.

5/01/97

CM/MW

Lowered boom for ECOR maintenance at 1720 GMT. Raised boom and RTS at 1915 GMT.

2/11/97

CM/MS

Raised boom and RTS at 2120 GMT.

2/10/97

MW/ZW

Lowered boom at 1950 GMT for ECOR maintenance. Unable to raise boom due to ECOR Sonic power problems. Left powered off at ground level overnight.

2/05/97

CAM/DC

Lowered 25M T/H at 1420 GMT for sensor checks. * R/R 25M humidity sensor. RTS at 2030 GMT. * Old - 231 * New - 234 (Cal 12/3/96)

2/04/97

Vendor

Tower Systems completed repairs and 25 M Temp/Hum RTS at 1845 GMT. Safety removed lock and tag out.

2/04/97

CAM/DC

Lowered 25M T/H at 2030 GMT for sensor checks. Raised boom and RTS at 2220 GMT.

2/03/97

Vendor

Tower Systems on site to make all necessary repairs to tower including replacing side B winch cable. System B booms lowered at 1530 GMT. Bottom limit switches were moved closer to base of rail and hard stops were removed. System B left down over night.

10/23/96

Safety

Elevator controls locked and tag out until tower systems can inspect and make repairs to the tower. 25 T/H still off line.

10/22/96

DC/CAM

Tower rails damaged during lowering. Unable to restore 25 T/H to operation until repairs can be made.

10/21/96

DC/CAM

Mentor on site. Lowered both 25 and 60 T/H for 6 month calibrations at 1800 GMT. Left at ground level overnight.

9/10/96

DKB/CAM

Lowered boom from 1942 to 2300 GMT for 60 MT maintenance.

8/28/96

TMG/GRS

T/S found the cable between the SC32A and bit driver bad. Replaced and tested OK at 1730 GMT.

8/26/96

GRS

SDS reports no files from new data logger.

8/23/96

CAM

* Replaced data logger and tested OK at 1902 GMT. * Old - 12415 * New - 6948

8/20/96

DKB/TMG

Performed mentors work request to T/S 21X problem. Found data logger bad.

8/15/96

GRS

SDS reported no files from 21X data logger and could not TIP. T/S found data logger bad. Informed mentor who is sending a replacement.

6/27/96

CAW/CAM

T/S comms problem continued at 2030 GMT. SDS still cannot link with 25 T/H. Switched back to bit driver power supply and tested OK.

6/21/96

CAM

T/S comms problem at 1800 GMT. Found loose fiber optic cable at the tower base. This did not correct the problem. Performed loop back test from the EFCOS and found it bad. R/R bit drivers and loop back tested OK from EFCOS using laptop. Still could not monitor. Receiving sequential comms interrupts. Changed from bit driver power supply to Eagan adapter cable power and established link from EFCOS. SDS down due to upgrade so further testing could not be performed at this time.

6/17/96

CAM

Sneakernet data files at mentors request at 2015MT. All data present. SDS still receiving short files. T/S continues.

6/6/96

TMG/CAM

T/S SDS receiving short data files. Mentor requested power cycle of data logger. Power cycled data logger at 1604 GMT.

5/21/96

CAW

Lowered tower for ECOR C1 installation at 1530 GMT.

4/3/96

CAM

T/S no comms with system since 4/2 2100 GMT. Reset bit drivers and EFCOS and tested OK at 1600 GMT.

4/2/96

TMG/CAM

Lowered tower at 1607 GMT for sensor calibration with mentor. Increased calibration multiplier to 1.4. RTS at 2240 GMT.

3/25/96

CAM/ZW

Lowered boom at 1855 GMT and * R/R humidity sensor on mentors request. RTS at 2145 GMT. * old - 226 * new - 231

3/19/96

TMG/CAM

Lowered sensor at 1905 GMT and tested for 1 hour. *R/R Hum sensor and tested overnight. RTS at 2130 GMT. * old - 234 * new - 226

 2/9/96

D. Cook

Installed instrument and tested OK.